OpenStudy (anonymous):

Which of these are the intercepts of the graph of y = 3x – 12? A. (4, 0), (0, -12) B. (12, 0), (0, 4) C. (3, 0), (0, 4) D. (4, 0), (0, 3) E. (-4, 0), (0, -12)

OpenStudy (anonymous):

For intercepts. Put y = 0 and find x intercept and put x = 0 to find y intercept so what will be the answer?

OpenStudy (anonymous):

(0,4) (0,-12)??

OpenStudy (anonymous):

(0,-12) is correct. but what will be (x,0)

OpenStudy (michele_laino):

please note that, I can write your equation, as below: \[\frac{ x }{ 4 }-\frac{ y }{ 12 }=1\]

OpenStudy (michele_laino):

or in better way: \[\frac{ x }{ 4 }+\frac{ y }{ -12 }=1\]

OpenStudy (anonymous):

it will be (4,0) (0,-12)

OpenStudy (anonymous):

Correct @jdmeduardo :)

OpenStudy (michele_laino):

that's right! @jdmeduardo